# Generalized quantum Zernike Hamiltonians: Polynomial Higgs-type algebras and algebraic derivation of the spectrum

### v1

@@ Abstract
^$\hat{\mathcal{H}}_{N}=\hat{p}_{1}^{2}+\hat{p}_{2}^{2}+\sum_{k=1}^{N}\gamma_{k}(\hat{q}_{1}\hat{p}_{1}+\hat{q}_{2}\hat{p}_{2})^{k},$
with canonical operators $\hat{q}_{i},\,\hat{p}_{i}$ and arbitrary coefficients $\gamma_{k}$. This two-dimensional quantum model, besides the conservation of the angular momentum, exhibits higher-order integrals of motion within the enveloping algebra of the Heisenberg algebra $\mathfrak{h}_{2}$. By constructing suitable combinations of these integrals, we uncover a polynomial Higgs-type symmetry algebra that, through an appropriate change of basis, gives rise to a deformed oscillator algebra. The associated structure function $\Phi$ is shown to factorize into two commuting components $\Phi=\Phi_{1}\Phi_{2}$. This framework enables an algebraic determination of the possible energy spectra of the model for the cases $N=2,3,4$, the case $N=1$ being canonically equivalent to the harmonic oscillator. Based on these findings, we propose two conjectures which generalize the results for all $N\geq 2$ and any value of the coefficients $\gamma_{k}$, that they are explicitly proven for $N=5$. In addition, all of these results can be interpreted as superintegrable perturbations of the original quantum Zernike system corresponding to $N=2$ which are also analyzed and applied to the isotropic oscillator on the sphere, hyperbolic and Euclidean spaces.
@@ 1 Introduction
The so-called Zernike system, emerged in the context of wavefront descriptions in optics [2], has turned out to be an extremely interesting model, with many physically and mathematically relevant applications. Besides being the first case in a hierarchy of (classically) superintegrable systems [3], its quantum version determines a new class of orthogonal polynomials, the Zernike polynomials, of great interest in the treatment of optical aberrations. In addition, it has been shown that the quantum symmetries of the system give rise to a cubic Higgs algebra $\mathfrak{sl}^{(3)}(2,\mathbb{R})$. The latter fact, in combination with recent work on polynomial algebras in connection with the superintegrability property (see e.g. [4] and references therein), suggest to look for a quantization of the generalized classical Zernike Hamiltonians considered in [3] in such a way that the corresponding quantum symmetry algebra is a higher-order polynomial Higgs-type algebra.
The two-dimensional Zernike system is defined by means of the following quantum Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ with eigenvalue equation (with normalized units $\hbar=1$)
^$\hat{\mathcal{H}}_{\rm Zk}\psi(\mathbf{q})=\left(-\nabla^{2}-\beta\,{\mathbf{q}}\boldsymbol{\cdot}\nabla-\alpha\bigl{(}{\mathbf{q}}\boldsymbol{\cdot}\nabla\bigr{)}^{2}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q}),$ (1.1)
@@
such that $\alpha$ and $\beta$ are two real parameters. It turns out that the Zernike system exhibits interesting structural features of great relevance in quantum optics, which have motivated its detailed study in recent years (see [5, 6, 7, 8, 9, 10] and references therein). Among these important properties, the Zernike system is superintegrable, that is, it admits two algebraically independent quantum observables that commute with the Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$. Therefore, the corresponding spectrum presents a maximal degeneracy. We also recall that, as shown in [2], among the possible values for the real parameters $\alpha$ and $\beta$, the case with $\alpha=-1$ and $\beta=-2$ is the only one that provides a self-adjoint operator ^$\hat{\mathcal{H}}_{\rm Zk}$ under the inner product over the unit disk with $|{\mathbf{q}}|\leq 1$, which has been fully studied by Pogosyan et al in [5], leading to the Zernike polynomials (see e.g. [11, 12] and references therein).
On the other hand, the classical Hamiltonian counterpart of ^$\hat{\mathcal{H}}_{\rm Zk}$ (1.1), also considered by Pogosyan et al in [13], adopts the following form
@@
where $q_{i}$ and $p_{i}$ are generic canonical variables fulfilling the Poisson bracket $\{q_{i},p_{j}\}=\delta_{ij}$. Obviously, the Hamiltonian $\mathcal{H}_{\rm Zk}$ determines a classical superintegrable system, meaning that it admits two constants of the motion, in this case quadratic in the momenta, which added to the Hamiltonian $\mathcal{H}_{\rm Zk}$ form a set of three functionally independent functions. It follows that all bounded trajectories are periodic, and for $\mathcal{H}_{\rm Zk}$ these are given by ellipses [13]. A first superintegrable generalization of the classical Zernike Hamiltonian (1.2) was considered and analyzed in [3] in the form
@@
where $\gamma_{k}$ are arbitrary coefficients that can be either pure imaginary or real numbers. Two constants of the motion for the Hamiltonian $\mathcal{H}_{N}$ are the usual angular momentum
$\mathcal{C}=q_{1}p_{2}-q_{2}p_{1}$ (1.4)
and a certain function $\mathcal{I}_{N}$, which is of $N^{{\rm th}}$-order in the momenta and has been obtained explicitly in [3] for arbitrary index $N$. The classical Zernike system $\mathcal{H}_{\rm Zk}$ is recovered from $\mathcal{H}_{N}$ (1.3) for the quadratic case with $N=2$ by setting
@@
In this context, $\mathcal{H}_{N}$ can be interpreted as a superintegrable perturbation of $\mathcal{H}_{\rm Zk}$ for $N\geq 3$, and such that the bounded trajectories are no longer ellipses, but some ‘deformation’ of them that increases with the values of $\gamma_{k}$ ($k\geq 3)$ as shown in [3]. More recently, an additional generalization of the Hamiltonian (1.3) has been proposed in [14] in terms of generic analytical functions $F\left(q_{1}p_{1}+q_{2}p_{2}\right)$, showing that the superintegrability property actually arises from the particular structure of the Hamiltonian.
The aim of this work is twofold. In the first place, we propose a generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ by imposing that the following requirements are satisfied:
@@
(ii) For $N=2$, the quantum Zernike system ^$\hat{\mathcal{H}}_{\rm Zk}$ in (1.1) is recovered.
(iii) The classical counterpart of ^$\hat{\mathcal{H}}_{N}$ is given by $\mathcal{H}_{N}$ in (1.3).
As second objective, we derive the corresponding spectrum of ^$\hat{\mathcal{H}}_{\rm Zk}$ by means of an algebraic procedure that involves the quantum symmetries of ^$\hat{\mathcal{H}}_{N}$, i.e., algebraically independent quantum observables commuting with ^$\hat{\mathcal{H}}_{N}$, which are shown to be of $N^{{\rm th}}$-order in the momenta.
Furthermore, we recall that the classical Zernike Hamiltonian $\mathcal{H}_{\rm Zk}$ (1.2) has been alternatively interpreted in [3] as the two-dimensional isotropic oscillator on the sphere, hyperbolic and Euclidean spaces (the former being just the Higgs oscillator), such that the parameter $\beta$ is related to the frequency, while $\alpha$ determines the constant curvature of the underlying space. As a direct consequence, for $N\geq 3$ our results provide superintegrable perturbations of such (curved) quantum oscillators that we also study in detail.
The structure of the paper is as follows: In section 2 we present the algebraic procedure used to obtain the quantum symmetries and spectrum for the generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ in (1.3). In section 3, our approach is first applied to the proper (quadratic) quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ (1.1), recovering known results. As a byproduct, the associated quantum curved oscillators are also deduced and studied. Sections 4 and 5 are devoted to the cubic $N=3$ and quartic $N=4$ systems, which are solved explicitly. In addition, we stress that the cubic case is shown to introduce a superintegrable perturbation of the quantum isotropic oscillator on the above three Riemannian spaces that are fully analyzed in section 4.2. Besides the two curved oscillators, this application also yields new quantum Euclidean oscillators which are named as either cubic spherical or hyperbolic perturbations.
Due to ordering problems in the quantization, the computational difficulties increase exponentially for higher values of $N$, so that general expression for arbitrary $N$ remains as an open problem; in fact, even in the classical framework [3], the general construction is quite involved. Nevertheless, we present in section 6 two conjectures for the spectrum of ^$\hat{\mathcal{H}}_{N}$ (1.3) with arbitrary $N$ along with the solution for $N=5$. The paper concludes with some remarks and some open problems whose detailed analysis will be considered in the future.
@@ 2 Generalized quantum Zernike Hamiltonians
$[\hat{q}_{i},\hat{q}_{j}]=[\hat{p}_{i},\hat{p}_{j}]=0,\qquad[\hat{q}_{i},\hat{p}_{j}]={\rm i}\,{\rm Id}\,\delta_{ij},\qquad\hat{q}_{i}\psi(\mathbf{q})=q_{i}\psi(\mathbf{q}),\qquad\hat{p}_{i}\psi(\mathbf{q})=-{\rm i}\,\frac{\partial\psi(\mathbf{q})}{\partial q_{i}},$ (2.1)
@@
^$\begin{split}\hat{\mathcal{H}}_{\rm Zk}&={\hat{p}}_{1}^{2}+{\hat{p}}_{2}^{2}-{\rm i}\beta({\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2})+\alpha({\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2})^{2}\\[2.0pt] &={\hat{\mathbf{p}}}^{2}-{\rm i}\beta({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})+\alpha({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})^{2}.\end{split}$ (2.2)
@@
^$\hat{\mathcal{H}}_{N}={\hat{\mathbf{p}}}^{2}+\sum_{k=1}^{N}\gamma_{k}({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})^{k},$ (2.3)
@@
^$\begin{split}\hat{\mathcal{H}}_{N}\psi(\mathbf{q})&=\left(-\nabla^{2}+\sum_{k=1}^{N}\gamma_{k}(-{\rm i})^{k}\bigl{(}{\mathbf{q}}\boldsymbol{\cdot}\nabla\bigr{)}^{k}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q})\\[4.0pt] &=\left(-\left(\frac{\partial^{2}}{\partial q_{1}^{2}}+\frac{\partial^{2}}{\partial q_{2}^{2}}\right)+\sum_{k=1}^{N}\gamma_{k}(-{\rm i})^{k}\left(q_{1}\frac{\partial}{\partial q_{1}}+q_{2}\frac{\partial}{\partial q_{2}~{}}\right)^{\!k}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q}),\end{split}$ (2.4)
which generalizes to arbitrary values of $N\geq 3$ the Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ (1.1), that is recovered with the choices (1.5). It is important to observe that, due to the term ordering problems arising in the quantization of $\mathcal{H}_{N}$ (1.3) as well as for the corresponding constants of the motion, there are many possibilities for the quantum construction. Nevertheless, we will show here that the Hamiltonian ^$\hat{\mathcal{H}}_{N}$ in (2.3) determines a quantum superintegrable system by explicitly finding its quantum symmetries and then computing its spectrum for arbitrary coefficients $\gamma_{k}$. As expected, for increasing values of $N$, some of the computations become extremely cumbersome, requiring the use of symbolic computer packages, such as the Wolfram Mathematica® software system.
The first step in our construction is to obtain the quantum symmetries of ^$\hat{\mathcal{H}}_{N}$ ensuring the superintegrability property. One of these symmetries clearly corresponds to the quantum angular momentum operator
@@
as it commutes with ${\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}}$. This shows that ^$\hat{\mathcal{H}}_{N}$ is a quantum integrable Hamiltonian. Due to the structure of the Hamiltonian, we expect the additional quantum symmetry for ^$\hat{\mathcal{H}}_{N}$ to be of higher-order in the momenta. Following this assumption, we consider the enveloping algebra of the Heisenberg algebra in two dimensions spanned by ${\hat{\mathbf{q}}}$ and ${\hat{\mathbf{p}}}$ (2.1), to which the identity operator ${\rm Id}$ is added as fifth generator, and introduce the following grading
@@
Next, we look for monomials $\hat{\mathcal{P}}_{\ell}^{(s)}({\hat{\mathbf{q}}},{\hat{\mathbf{p}}})$ of degree $\ell\leq 2N$ such that the total grading of the monomial is equal to 0. The grading is assigned via the prescription
$|{\hat{q}}_{1}^{a}\,{\hat{q}}_{2}^{b}\,{\hat{p}}_{1}^{c}\,{\hat{p}}_{2}^{d}|=a+b-c-d.$
In this way, we construct the following operators of $N^{{\rm th}}$-order in the momenta
${\hat{I}}_{N}={\hat{p}}_{2}^{2}+\sum_{\ell}^{2N}\sum_{s}^{{\rm dim}\hat{\mathcal{P}}_{\!\ell}}A_{\ell}^{(s)}\hat{\mathcal{P}}_{\ell}^{(s)},\qquad{\hat{I}}_{N}^{\prime}={\hat{p}}_{1}^{2}+\sum_{\ell}^{2N}\sum_{s}^{{\rm dim}\hat{\mathcal{P}}_{\!\ell}}B_{\ell}^{(s)}\hat{\mathcal{P}}_{\ell}^{(s)},$ (2.6)
where $A_{\ell}^{(s)}$ and $B_{\ell}^{(s)}$ are certain coefficients to be determined. These operators provide quantum symmetries for the Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.3) by imposing the condition
@@
which determines the explicit form for the monomials $\hat{\mathcal{P}}_{\ell}^{(s)}$ together with the values of $A_{\ell}^{(s)}$ and $B_{\ell}^{(s)}$. Provided that these quantum symmetries have been found, it follows that the sets (^$(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N})$ and (^$(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}^{\prime}_{N})$ are formed each by three algebraically independent operators, so that ^$\hat{\mathcal{H}}_{N}$ is a quantum superintegrable Hamiltonian. As happens in the classical case (see [3]), the four operators (^$(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N},{\hat{I}}^{\prime}_{N})$ satisfy an algebraic dependence relation.

### v2

@@ Abstract
$\hat{\mathcal{H}}_{N}=\hat{p}_{1}^{2}+\hat{p}_{2}^{2}+\sum_{k=1}^{N}\gamma_{k}(\hat{q}_{1}\hat{p}_{1}+\hat{q}_{2}\hat{p}_{2})^{k},$
with canonical operators $\hat{q}_{i},\,\hat{p}_{i}$ and arbitrary coefficients $\gamma_{k}$. This two-dimensional quantum model, besides the conservation of the angular momentum, exhibits higher-order integrals of motion within the enveloping algebra of the Heisenberg algebra in two dimensions. By constructing suitable combinations of these integrals, we uncover a polynomial Higgs-type symmetry algebra that, through an appropriate change of basis, gives rise to a deformed oscillator algebra. The associated structure function $\Phi$ is shown to factorize into two commuting components $\Phi=\Phi_{1}\Phi_{2}$. This framework enables an algebraic determination of the possible energy spectra of the model for the cases $1\leq N\leq 5$, the case $N=1$ being canonically equivalent to the harmonic oscillator. Based on these findings, we propose two conjectures which generalize the results for all $N\geq 1$ and any value of the coefficients $\gamma_{k}$. In addition, all of these results can be interpreted as higher-order superintegrable perturbations of the original quantum Zernike system corresponding to $N=2$, which are also analyzed and applied to the isotropic oscillator on the sphere, hyperbolic and Euclidean spaces.
@@ 1 Introduction
The so-called Zernike system, that emerged in the context of wavefront descriptions in optics [3], has turned out to be an extremely interesting model, with many physically and mathematically relevant applications. Besides being the first case in a hierarchy of (classically) superintegrable systems [4], its quantum version determines a new class of orthogonal polynomials, the Zernike polynomials, of great interest in the treatment of optical aberrations. In addition, it has been shown that the quantum symmetries of the system give rise to a cubic Higgs algebra $\mathfrak{sl}^{(3)}(2,\mathbb{R})$. The latter fact, in combination with recent work on polynomial algebras in connection with the superintegrability property (see e.g. [5] and references therein), suggest to look for a quantization of the generalized classical Zernike Hamiltonians considered in [4] in such a way that the corresponding quantum symmetry algebra is a higher-order polynomial Higgs-type algebra.
The two-dimensional Zernike system is defined by means of the following quantum Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$ with eigenvalue equation (with normalized units $\hbar=1$)
$\hat{\mathcal{H}}_{\rm Zk}\psi(\mathbf{q})=\left(-\nabla^{2}-\beta\,{\mathbf{q}}\boldsymbol{\cdot}\nabla-\alpha\bigl({\mathbf{q}}\boldsymbol{\cdot}\nabla\bigr)^{2}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q}),$ (1.1)
@@
such that $\alpha$ and $\beta$ are two real parameters. It turns out that the Zernike system exhibits interesting structural features of great relevance in quantum optics, which have motivated its detailed study in recent years (see [6, 7, 8, 9, 10, 11] and references therein). Among these important properties, the Zernike system is superintegrable, that is, it admits two algebraically independent quantum observables that commute with the Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$. Therefore, the corresponding spectrum presents a maximal degeneracy. We also recall that, as shown in [3], among the possible values for the real parameters $\alpha$ and $\beta$, the case with $\alpha=-1$ and $\beta=-2$ is the only one that provides a self-adjoint operator $\hat{\mathcal{H}}_{\rm Zk}$ under the inner product over the unit disk with $|{\mathbf{q}}|\leq 1$, which has been fully studied by Pogosyan et al in [6], leading to the Zernike polynomials (see e.g. [12, 13, 14] and references therein).
On the other hand, the classical Hamiltonian counterpart of $\hat{\mathcal{H}}_{\rm Zk}$ (1.1), also considered by Pogosyan et al in [15], adopts the following form
@@
where $q_{i}$ and $p_{i}$ are generic canonical variables fulfilling the Poisson bracket $\{q_{i},p_{j}\}=\delta_{ij}$. The Hamiltonian $\mathcal{H}_{\rm Zk}$ determines a classical superintegrable system (in the Liouville sense [16]), which means that it admits two constants of the motion, in this case quadratic in the momenta. Explicitly, these are the usual angular momentum
$\mathcal{C}=q_{1}p_{2}-q_{2}p_{1}$ (1.3)
together with a function given by
$\mathcal{I}=p_{2}^{2}-{\rm i}\beta q_{2}p_{2}+\alpha\bigl(q_{1}^{2}+q_{2}^{2}\bigr)p_{2}^{2}.$ (1.4)
Thus, the Hamiltonian $\mathcal{H}_{\rm Zk}$ Poisson-commutes with $\mathcal{C}$ and $\mathcal{I}$, although these constants are not in involution:
$\big\{\mathcal{H}_{\rm Zk},\mathcal{C}\big\}=\big\{\mathcal{H}_{\rm Zk},\mathcal{I}\}=0,\qquad\big\{\mathcal{C},\mathcal{I}\}\neq 0.$
In this respect, we recall that the existence of one constant of the motion implies that $\mathcal{H}_{\rm Zk}$ determines a classical integrable system. Moreover, as the three functions $\mathcal{H}_{\rm Zk}$, $\mathcal{C}$ and $\mathcal{I}$ are functionally independent, the classical Zernike Hamiltonian is superintegrable, and therefore endowed with the maximum number of constants of the motion for this dimension. As a consequence, all bounded trajectories are closed and periodic, which for $\mathcal{H}_{\rm Zk}$ are given by ellipses [15].
In addition, another constant of the motion can be obtained immediately from $\mathcal{I}$ (1.4) in the form
$\mathcal{I}^{\prime}=p_{1}^{2}-{\rm i}\beta q_{1}p_{1}+\alpha\bigl(q_{1}^{2}+q_{2}^{2}\bigr)p_{1}^{2},\qquad\big\{\mathcal{H}_{\rm Zk},\mathcal{I}^{\prime}\}=0,$ (1.5)
which cannot be functionally independent with respect to the previous ones, thereby verifying the following relation
$\mathcal{H}_{\rm Zk}=\mathcal{I}+\mathcal{I}^{\prime}-\alpha\mathcal{C}^{2}.$
The introduction of $\mathcal{I}^{\prime}$ enables the construction of the symmetry algebra of $\mathcal{H}_{\rm Zk}$, understood as the algebra closed by its constants of the motion. According to [15], if we define
$\begin{split}\mathcal{L}_{1}&:=\frac{1}{2}\mathcal{C},\qquad\mathcal{L}_{2}:=\frac{1}{2}\bigl(\mathcal{I}^{\prime}-\mathcal{I}\bigr),\\[2.0pt] \mathcal{L}_{3}&:=\{\mathcal{L}_{1},\mathcal{L}_{2}\}=\left(1+\alpha\bigl(q_{1}^{2}+q_{2}^{2}\bigr)\right)p_{1}p_{2}-\frac{1}{2}{\rm i}\beta(q_{1}p_{2}+q_{2}p_{1}),\end{split}$ (1.6)
we obtain the following Poisson brackets
$\{\mathcal{L}_{1},\mathcal{L}_{2}\}=\mathcal{L}_{3},\qquad\{\mathcal{L}_{1},\mathcal{L}_{3}\}=-\mathcal{L}_{2},\qquad\{\mathcal{L}_{2},\mathcal{L}_{3}\}=\big(\beta^{2}-2\alpha\mathcal{H}_{\rm Zk}\big)\mathcal{L}_{1}-8\alpha^{2}\mathcal{L}_{1}^{3}.$ (1.7)
This structure determines a cubic polynomial Poisson algebra, which for the values $\alpha=-1$ and $\beta=-2$ [15] reproduces the so-called Higgs algebra, i.e., the symmetry algebra of the Higgs oscillator on the sphere [17]. In fact, we recall that if different values of the parameter $\alpha$ are allowed, the classical Zernike Hamiltonian $\mathcal{H}_{\rm Zk}$ (1.2) has been alternatively interpreted in [4] as the two-dimensional isotropic oscillator on the sphere, hyperbolic and Euclidean spaces for $\alpha<0$, $\alpha>0$ and $\alpha=0$, respectively. Therefore, from this point of view, the parameter $\alpha$ determines the constant curvature of the underlying space, while $\beta$ is related to the frequency of the corresponding oscillator. Observe that for the Euclidean oscillator with $\alpha=0$, the polynomial algebra (1.7) reduces to the Lie algebra $\mathfrak{so}(3)$. It should also be noted that, very recently, it has been shown in [18] that the ‘imaginary $\beta$-term’ in $\mathcal{H}_{\rm Zk}$ (1.2) is provided by a vector potential that determines a magnetic field, which is a pure gauge and can be eliminated by a canonical transformation for any value of $\alpha$.
A first superintegrable generalization of the classical Zernike Hamiltonian (1.2) was considered and analyzed in [4] in the form
@@
where $\gamma_{k}$ are arbitrary coefficients that can be either pure imaginary or real numbers. Two constants of the motion for the Hamiltonian $\mathcal{H}_{N}$ are the angular momentum (1.3) a certain function $\mathcal{I}_{N}$, which is of $N^{{\rm th}}$-order in the momenta and has been obtained explicitly in [4] for arbitrary index $N$. The classical Zernike system $\mathcal{H}_{\rm Zk}$ is recovered from $\mathcal{H}_{N}$ (1.8) for the quadratic case with $N=2$ by setting
@@
In this context, $\mathcal{H}_{N}$ can be interpreted as a superintegrable perturbation of $\mathcal{H}_{\rm Zk}$ for $N\geq 3$, and such that the bounded trajectories, albeit preserving the property of being closed and periodic, are no longer ellipses, but some ‘deformation’ of them that increases with the values of $\gamma_{k}$ ($k\geq 3)$, as shown in [4]. More recently, an additional generalization of the Hamiltonian (1.8) has been proposed in [19] in terms of generic analytical functions $F\left(q_{1}p_{1}+q_{2}p_{2}\right)$, showing that the superintegrability property actually arises from the particular structure of the Hamiltonian.
The aim of this work is twofold. In the first place, we propose a generalized quantum Zernike Hamiltonian $\hat{\mathcal{H}}_{N}$ by imposing that the following requirements are satisfied:
@@
(ii) For $N=2$, the quantum Zernike system $\hat{\mathcal{H}}_{\rm Zk}$ in (1.1) is recovered.
(iii) The classical counterpart of $\hat{\mathcal{H}}_{N}$ is given by $\mathcal{H}_{N}$ in (1.8).
As second objective, we derive the corresponding spectrum of $\hat{\mathcal{H}}_{\rm Zk}$ by means of an algebraic procedure that involves the quantum symmetries of $\hat{\mathcal{H}}_{N}$, i.e., algebraically independent quantum observables commuting with $\hat{\mathcal{H}}_{N}$, which are shown to be of $N^{{\rm th}}$-order in the momenta.
Furthermore, similarly to what occurs with $\mathcal{H}_{\rm Zk}\equiv\mathcal{H}_{2}$, the quantum Zernike system $\hat{\mathcal{H}}_{\rm Zk}\equiv\hat{\mathcal{H}}_{2}$ can be interpreted as a quantum oscillator on the sphere, hyperbolic and Euclidean spaces; the latter actually corresponds to the case with $N=1$ ($\gamma_{2}=\alpha=0$). As a direct consequence of this result, our findings provide superintegrable perturbations of such (curved) quantum oscillators for $N\geq 3$, which we also study in detail.
The structure of the paper is as follows: In section 2 we present the algebraic procedure used to obtain the quantum symmetries and spectrum for the generalized quantum Zernike Hamiltonian $\hat{\mathcal{H}}_{N}$ in (1.8). In section 3, our approach is first applied to the proper (quadratic) quantum Zernike Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$ (1.1), recovering known results. As a byproduct, the associated quantum curved oscillators are also deduced and studied. Section 4 is devoted to the cubic system, with $N=3$, which is explicitly solved. In addition, we stress that the cubic case is shown to introduce a superintegrable perturbation of the quantum isotropic oscillator on the above three Riemannian spaces that are fully analyzed in section 4.2. In particular, two types of perturbations arise, called spherical or hyperbolic cubic perturbations, which hold for each of the quantum oscillators in the sphere, hyperbolic and Euclidean spaces. Therefore, in addition to the two curved oscillators, which cover the Higgs oscillator, this application also yields two new Euclidean quantum oscillators.
Due to ordering problems in the quantization, the computational difficulties increase exponentially for higher values of $N$. To facilitate the reading of the manuscript, the fourth-order system with $N=4$ is addressed in Appendix A, where the construction is developed along the same lines as for the previous cubic system and the corresponding results are presented in detail. The main results for the fifth-order system with $N=5$ are summarized in Appendix B.
It should be observed that the general expression for arbitrary $N$ remains as an open problem; in fact, even in the classical framework [4], the general construction is quite involved. Nevertheless, taking into account the results up to $N=5$, in section 5 we present two conjectures for the spectrum of $\hat{\mathcal{H}}_{N}$ (1.8) with arbitrary $N$. The paper concludes with some remarks and some open problems whose detailed analysis will be considered in the future.
@@ 2 Generalized quantum Zernike Hamiltonians
$\hat{q}_{i},\hat{q}_{j}]=[\hat{p}_{i},\hat{p}_{j}]=0,\qquad[\hat{q}_{i},\hat{p}_{j}]={\rm i}\,{\rm Id}\,\delta_{ij},\qquad\hat{q}_{i}\psi(\mathbf{q})=q_{i}\psi(\mathbf{q}),\qquad\hat{p}_{i}\psi(\mathbf{q})=-{\rm i}\,\frac{\partial\psi(\mathbf{q})}{\partial q_{i}},$ (2.1)
@@
$\begin{split}\hat{\mathcal{H}}_{\rm Zk}&={\hat{p}}_{1}^{2}+{\hat{p}}_{2}^{2}-{\rm i}\beta({\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2})+\alpha({\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2})^{2}\\[2.0pt] &={\hat{\mathbf{p}}}^{2}-{\rm i}\beta({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})+\alpha({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})^{2}.\end{split}$ (2.2)
@@
$\hat{\mathcal{H}}_{N}={\hat{\mathbf{p}}}^{2}+\sum_{k=1}^{N}\gamma_{k}({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})^{k},$ (2.3)
@@
$\begin{split}\hat{\mathcal{H}}_{N}\psi(\mathbf{q})&=\left(-\nabla^{2}+\sum_{k=1}^{N}\gamma_{k}(-{\rm i})^{k}\bigl({\mathbf{q}}\boldsymbol{\cdot}\nabla\bigr)^{k}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q})\\[4.0pt] &=\left(-\left(\frac{\partial^{2}}{\partial q_{1}^{2}}+\frac{\partial^{2}}{\partial q_{2}^{2}}\right)+\sum_{k=1}^{N}\gamma_{k}(-{\rm i})^{k}\left(q_{1}\frac{\partial}{\partial q_{1}}+q_{2}\frac{\partial}{\partial q_{2}~}\right)^{\!k}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q}),\end{split}$ (2.4)
which generalizes to arbitrary values of $N\geq 3$ the Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$ (1.1), that is recovered with the choices (1.9). It is important to observe that, due to the term ordering problems arising in the quantization of $\mathcal{H}_{N}$ (1.8), as well as for the corresponding constants of the motion, there are many possibilities for the quantum construction. Nevertheless, we will show here that the Hamiltonian $\hat{\mathcal{H}}_{N}$ in (2.3) determines a quantum superintegrable system by explicitly finding its quantum symmetries and then computing its spectrum for arbitrary coefficients $\gamma_{k}$. As expected, for increasing values of $N$, some of the computations become extremely cumbersome, requiring the use of symbolic computer packages, such as the Wolfram Mathematica® software system.
The first step in our construction is to obtain the quantum symmetries of $\hat{\mathcal{H}}_{N}$ ensuring the superintegrability property. One of these symmetries corresponds to the quantum angular momentum operator
@@
as it commutes with $^{\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}}$. This shows that $\hat{\mathcal{H}}_{N}$ is a quantum integrable Hamiltonian. Due to the structure of the Hamiltonian, we expect the additional quantum symmetry for $\hat{\mathcal{H}}_{N}$ to be of higher-order in the momenta. Following this assumption, we consider the enveloping algebra of the Heisenberg algebra in two dimensions spanned by $^{\hat{\mathbf{q}}}$ and $^{\hat{\mathbf{p}}}$ (2.1), to which the identity operator $\rm Id$} is added as fifth generator, and introduce the following grading
@@
[…]
$|{\hat{q}}_{1}^{a}\,{\hat{q}}_{2}^{b}\,{\hat{p}}_{1}^{c}\,{\hat{p}}_{2}^{d}|=a+b-c-d,$
where $a$, $b$, $c$ and $d$ are integers. In this way, we construct the following operators of $N^{{\rm th}}$-order in the momenta
${\hat{I}}_{N}={\hat{p}}_{2}^{2}+\sum_{{\ell=1}}^{2N}\sum_{{s=1}}^{{\rm dim}\hat{\mathcal{P}}_{\!\ell}}A_{\ell}^{(s)}\hat{\mathcal{P}}_{\ell}^{(s)},\qquad{\hat{I}}_{N}^{\prime}={\hat{p}}_{1}^{2}+\sum_{{\ell=1}}^{2N}\sum_{{s=1}}^{{\rm dim}\hat{\mathcal{P}}_{\!\ell}}B_{\ell}^{(s)}\hat{\mathcal{P}}_{\ell}^{(s)},$ (2.6)
where $A_{\ell}^{(s)}$ and $B_{\ell}^{(s)}$ are certain coefficients to be determined. These operators provide quantum symmetries for the Hamiltonian $\hat{\mathcal{H}}_{N}$ (2.3) by imposing the condition
@@
which determines the explicit form for the monomials $\hat{\mathcal{P}}_{\ell}^{(s)}$ together with the values of $A_{\ell}^{(s)}$ and $B_{\ell}^{(s)}$. Provided that these quantum symmetries have been found, it follows that the sets $(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N})$ and $(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}^{\prime}_{N})$ are formed each by three algebraically independent operators, so that $\hat{\mathcal{H}}_{N}$ is a quantum superintegrable Hamiltonian. As happens in the classical case (see [4]), the four operators $(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N},{\hat{I}}^{\prime}_{N})$ satisfy an algebraic dependence relation.